OpenStudy (anonymous):

Solve the quadratic equation by completing the sqaure. x^2+12x+13=0

OpenStudy (vijay):

-6 - sqrt23 and sqrt (23)- 6

OpenStudy (anonymous):

\[x^2+12x+13=0\] \[x^2+12x=-13\] \[x^2+12x+36=-13+36\] \[(x+6)^2=23\] \[x+6=\pm \sqrt{23}\] \[x=-6\pm \sqrt{23}\]